# High Gain Fusion Target Design using Generative Artificial Intelligence

### v5

@@ Abstract
By returning to the topological basics of fusion target design, Generative Artificial Intelligence (genAI) is used to specify how to initially configure and drive the optimally entangled topological state, and stabilize that topological state from disruption. This can be applied to all methods; including tokamaks, laser-driven schemes, and pulsed-power driven schemes. This target design philosophy has the potential to yield practical, room temperature targets that could yield up to 10 GJ of energy, driven by as little as 3 MJ of absorbed energy. The genAI is based on the concept of Ubuntu that replaces the Deep Convolutional Neural Network approximation of a functional, with the formula for the generating functional of a canonical transformation from the domain of the canonical field momentums and fields, to the domain of the canonical momentums and coordinates, that is the Reduced Order Model. This formula is a logical process of renormalization, that has the potential to enable Heisenberg’s canonical approach to field theory, via calculation of the S-matrix, given observation of the fields. This can be viewed as topological characterization and control of collective, that is complex, systems.
@@ I Return to Topological Basics
MagLIF has a metastable, yet modestly inefficient, 2D implosion of the DT igniter and efficient 2D burn propagation. Laser ICF has a highly-unstable 3D implosion of the DT igniter and an inefficient 3D burn. Because of this both can barely ignite and burn DT ice. Ubuntu Fusion Target Technology has a metastable highly-efficient 3D implosion of the DT igniter and modestly efficient 2.5D burn, so that it can ignite and burn a “solid fuel” (i.e., a fuel that is a solid at room temperature and harder to ignite and burn than DT ice, but may have greater yield).
@@ II The Ubuntu Fusion Target
For brevity, we will describe the laser targets and drives. The Ubuntu Fusion Target is a simple, easy to fabricate, room temperature target. This is in contrast to the impractical and complicated MagLIF DT ice burner, and current ICF target designs. As shown in Fig. 3, the target consists of a cylinder of beryllium, coated on the inside with solid fuel, and filled with DT gas. The target is driven by four lasers whose illumination patterns on the cylinder are shaped with phase plates, as shown in Fig. 4, either to form a Z-pinch or an RFP. The pattern of the laser drive effectively makes the target plasma into a substantial MIFED coil. The implosion is not driven by radiation pressure nor ablation pressure, but by magnetic pressure which increases as the target is imploded. Unlike pulsed-power, which must physically touch the plasma, the lasers stand off from the target and only require four small lines of sight in a plane perpendicular to the axial direction. Ubuntu genAI will be used to optimize the target and drive design, including the axial structure. It will also be used to ponderomotively stabilize the metastable stagnation.
@@
Existing simulations and experiments indicate that the fuel can be assembled, ignition conditions might well be obtained in the igniter, and conditions can be obtained for a reasonably efficient 2.5D burn of the fuel; using as little as 3 MJ of energy on target. The yield could be up to 10 GJ.
@@ Acknowledgements.
Thanks is given to CSIRO for supporting the early parts of this research through their Science Leaders Program, and the Institute des Hautes Etudes Scientifique (IHES) for hosting a stay where the initial details of this theory were formalized. Thanks is also given to both Stéphane Mallat and Joan Bruna for many useful discussions, and communication of many of their mathematical results before they were presented or published. The mentorship of both Ted Frankel and Michael Freedman in topology and the geometry of physics is the foundation on which this work was constructed. Finally, thanks is given to the University of Western Australia, John Hedditch, and Ian MacArthur for their help in understanding many of the finer points of Quantum Field Theory. The subject matter of this Letter is the basis of the US Patent Application US20250116976-A1, entitled ”Systems and Methods for Controlling Complex Systems”, and the Provisional US Patent Application 63/760917, entitled ”Inertial Fusion Energy Generation Systems and Methods”.

### v6

@@ Abstract
By returning to the topological basics of fusion target design, Generative Artificial Intelligence (genAI) is used to specify how to initially configure and drive the optimally entangled topological state, and stabilize that topological state from disruption. This can be applied to all methods; including tokamaks, laser-driven schemes, and pulsed-power driven schemes. The result is practical, room temperature targets that can yield up to 10 GJ of energy, driven by as little as 3 MJ of absorbed energy. The genAI is based on the concept of Ubuntu that replaces the Deep Convolutional Neural Network approximation of a functional, with the formula for the generating functional of a canonical transformation from the domain of the canonical field momentums and fields, to the domain of the canonical momentums and coordinates, that is the Reduced Order Model. This formula is a logical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix, given observation of the fields. This can be viewed as topological characterization and control of collective, that is complex, systems.
@@ I Return to Topological Basics
MagLIF has a metastable, yet modestly inefficient, 2D implosion of the DT igniter and efficient 2D burn propagation. Laser ICF has a highly-unstable 3D implosion of the DT igniter and an inefficient 3D burn. Because of this both can barely ignite and burn DT ice. Ubuntu Fusion Target Technology has a metastable highly-efficient 3D implosion of the DT igniter and modestly efficient 2.5D burn, so that it can ignite and burn a “solid fuel” (i.e., like $^{6}\text{LiD}$ or $^{11}\text{BH}$, a fuel that is a solid at room temperature and harder to ignite and burn than DT ice, but may have greater yield).
@@ II The Ubuntu Fusion Target
For brevity, we will describe the laser targets and drives. The Ubuntu Fusion Target is a simple, easy to fabricate, room temperature target. This is in contrast to the impractical and complicated MagLIF DT ice burner, and current ICF target designs. As shown in Fig. 3, the target consists of a cylinder of beryllium, coated on the inside with solid fuel, and filled with DT gas. The target is driven by four lasers whose illumination patterns on the cylinder are shaped with phase plates, as shown in Fig. 4, either to form a Z-pinch or an RFP. The pattern of the laser drive effectively makes the target plasma into a substantial laser-driven Magneto-Inertial Fusion Electrical Discharge System (MIFEDS) coil [16]. The implosion is not driven by radiation pressure nor ablation pressure, but by magnetic pressure which increases as the target is imploded. Unlike pulsed-power, which must physically touch the plasma, the lasers stand off from the target and only require four small lines of sight in a plane perpendicular to the axial direction. Ubuntu genAI will be used to optimize the target and drive design, including the axial structure. It will also be used to ponderomotively stabilize the metastable stagnation.
@@
Existing simulations and experiments indicate that the fuel can be assembled, ignition conditions can be obtained in the igniter, and conditions can be obtained for an efficient 2D burn of the fuel; using as little as 3 MJ of energy on target. The yield can be up to 10 GJ.
@@ Acknowledgements.
Thanks is given to CSIRO for supporting the early parts of this research through their Science Leaders Program, and the Institute des Hautes Etudes Scientifique (IHES) for hosting a stay where the initial details of this theory were formalized. Thanks is also given to both Stéphane Mallat and Joan Bruna for many useful discussions, and communication of many of their mathematical results before they were presented or published. The mentorship of both Ted Frankel and Michael Freedman in topology and the geometry of physics is the foundation on which this work was constructed. Finally, thanks is given to the University of Western Australia, John Hedditch, and Ian MacArthur for their help in understanding many of the finer points of Quantum Field Theory.